Kinetic Energy- KE-1 a) A 1,500kg car is traveling at 30.0m/s. How many joules of kinetic enorekye energy as the KE-2 a) Two objects with masses m -20.0 have? does the car b) Find the speed in km/hr for a 7.00kg bowling ball to have the same etic energy Find the ratio E/Ex g and my-60.0kg are both traveling with the same speed V. b) Two objects with identical masses m are traveling with speeds V1 20.0ms c) In part a), my 3mi. In part b) V-3V,. Why are the two ratios of Eo/Ext not the same? Find the ratio Ev/Ex KE-3 Find the kinetic energy of a 65.0kg runner who can travel 1,500m in 4 minutes.
Answer #1

here,

1)

mass of car , m1 = 1500 kg

speed of car , u1 = 30 m/s

the kinetic energy of the car , KE = 0.5 * m1 * u1^2

KE = 0.5 * 1500 * 30^2 J

KE = 6.75 * 10^5 J

b)

mass of bowling ball , m2 = 7 kg

let the speed of bowling ball be u2

as they have same kinetic energy

KE = 0.5 *m2 * u2^2

6.75 * 10^5 = 0.5 * 7 * u2^2

solving for u2

u2 = 439.2 m/s

the speed of second ball is 439.2 m/s